OpenStudy (anonymous):

GEOMETRY HELP?(: Sabrina is fencing her yard, which is shaped as a rhombus. She used (6x – 7) meters of fencing material for one side and (3x+ 5) meters for the adjacent side. What is the perimeter of the yard?

OpenStudy (anonymous):

72 meters 68 meters 116 meters 92 meters

OpenStudy (anonymous):

I got 72 but I'm not sure

OpenStudy (anonymous):

(6x-7)=(3x+5) add 7 &5 you get 12 6x=3x+12 subtract 3 3x=12 3 goes into 12 4 times so x=4 put the 4 in place of the x ; 6(4)-7=17 3(4)+5=17 that makes each of the sides 17, & 4 * 17 equals what.?

OpenStudy (anonymous):

I got everything right except when I added 34 and 34 I somehow ended up with 72. Thanks a lot man(:

OpenStudy (anonymous):

No prob